Unipart, a manufacturer of auto parts, is considering two B2B (Business-to-Business) marketplaces to purchase its MRO (MainteHI I NEED TO DO A DECISION TREE OF THIS, PLEASE HELP!
Answer #1

Answer :

Parts4u.com

AllMRO.com's

prices

same

same

service levels

same

same

lead times

same

same

fee structures

5 percent commission

1 percent commission

subscription fee

Nil

$10 million

Unipart spends about $150 million on MRO

first year $150*10% = 15

so first year expected to spend 135 million

second year another 10% drop = 150*10+10 = 30

so second year expected to spend 120 million

And Unipart uses a discount rate of 20 percent

By keeping all this Unipart is spending 120 to 150 million per year. So here we compare this two B2B by taking concept of Time value of money. The time value of money concept says that “The value of Dollar today is more worthy then tomorrow” so by observing AllMRO.com's is taking subscription fee of $10 million as Parts4u.com not taking any subscription fee so as per my view we have to select Parts4u.com.